OpenStudy (anonymous):

Write a recursive formula for the sequence 15, 26, 48, 92, 180, ... Then find the next term. A) an= 2an – 1 – 4, where a1 = 15; 356 B) an= 2an – 4, where a1 = 15; 356 C) an= 4 + 11 · 2n–1, where a1 = 15; 356 D) an= 3an – 1 – 19, where a1 = 15; 356

OpenStudy (immings):

2(an)-4 2(15)-4=26 2(26)-4=48 2(48)-4=92 2(92)-4=180 2(180)-4=356

OpenStudy (anonymous):

Sorry I was eating dinner um if I can ask how exactly did you choose B? What makes it recursive?